# Brain-Inspired Stepwise Patch Merging for Vision Transformers

### v1

@@ Introduction
Setting. Object detection and instance segmentation experiments were conducted on the challenging COCO benchmark (Lin et al. 2014). All models were trained on the training set comprising 118k images and evaluated on the validation set with 5k images. We validated the effectiveness of different backbones using Mask R-CNN (He et al. 2017). Before training, the weights pre-trained on ImageNet-1K were used to initialize the backbone, and the newly added layers were initialized using the Xavier initialization method (Glorot and Bengio 2010). Our models were trained with a batch size of 16 on 8 NVIDIA A100 GPUs and optimized using the AdamW optimizer (Loshchilov and Hutter 2017) with an initial learning rate of $1\times 10^{-4}$.
@@
Result. As shown in Tab. 1, incorporating the SPM framework into the PVT resulted in significant improvements of 4.1%, 2.6%, and 1.3% in the Tiny, Small, and Medium models, respectively, for the object detection task. Notably, the SPM framework also demonstrated substantial improvements in the instance segmentation task. Several observations can be made by analyzing the detection results of models with different sizes. Models integrated with SPM show marked improvements in detecting medium-sized and large objects. This enhancement is attributed to the original patch merging’s relatively singular and small receptive fields, whereas the MSA module integrates features with diverse receptive fields, enabling the model to more accurately capture long-range dependencies. Moreover, there is a significant improvement in detecting small objects. Although larger models are typically better at modeling global relationships, the disruption of local information may hinder small objects from establishing complete semantic information, leading to missed detections. The GLE module addresses this by enhancing the perception of local discriminative information, resulting in consistent improvements in the detection performance of small objects across models of different sizes.
Semantic Segmentation on ADE20K
@@
ResNet18 (He et al. 2016) 15.5 32.9
PVT-Tiny (Wang et al. 2021) 17.0 35.7
PVT-Tiny (HAFA) (Chen et al. 2023) 18.7 40.1
@@
ResNet50 (He et al. 2016) 28.5 36.7
PVT-Small (Wang et al. 2021) 28.2 39.8
PVT-Small (HAFA) (Chen et al. 2023) 29.9 43.8
@@
ResNet101 (He et al. 2016) 47.5 38.8
ResNeXt101-32$\times$4d (Xie et al. 2017) 47.1 39.7
PVT-Medium (Wang et al. 2021) 48.0 41.6
@@
Figure 4: Visualization of the attention map includes the original images (the first column), the visualization of the attention map for each head after using SPM (the second and third columns), and without using SPM (the fourth and fifth columns). The red five-pointed star in the original images represents the position of the query.
@@
ICCV 2021 Swin-T (Liu et al. 2021) 29.0 81.3
Swin-T (HAFA) (Liu et al. 2021) 29.1 81.7
@@
CVPR 2023 NAT-Mini (Hassani et al. 2023) 20.0 81.8
@@
Setting. The ADE20K dataset (Zhou et al. 2017) is a widely utilized benchmark for semantic segmentation, comprising 150 categories with 20,210 images for training, 2,000 images for validation, and 3,352 images for testing. All the methods compared were evaluated using the Semantic FPN framework (Kirillov et al. 2019). The backbone network of our method was initialized with the pre-trained ImageNet-1k model, and the newly added layers were initialized using the Xavier initialization method. The initial learning rate was set to 0.0001, and the model was optimized using the AdamW optimizer. We trained our models for 40,000 iterations with a batch size of 16 on eight NVIDIA A100 GPUs. The learning rate followed a polynomial decay schedule with a power of 0.9. During training, images were randomly resized and cropped to $512\times 512$ pixels. For testing, images were rescaled to have a shorter side of 512 pixels.
Figure 5: Visualization comparison of the Effective Receptive Field of PVT-Tiny before (the first row) and after applying our SPM (the second row), using the output from the last three stages. Each ERF image is generated by averaging over 5000 $256\times 256$ images from ImageNet-1K validation set.
Result. As shown in Tab. 3, the integration of the SPM framework led to a significant enhancement in the semantic segmentation task. Specifically, the performance of the Tiny, Small, and Large models improved by 5.8%, 6.1%, and 3.7%, respectively. It is evident that the improvement achieved by SPM in segmentation tasks surpasses that in classification and detection tasks. Interestingly, in both classification and detection tasks, the relative improvement brought by SPM compared to the base model gradually decreases as the model size increases. However, on the ADE20K dataset, the performance gain of PVT-Small exceeds that of PVT-Tiny, with improvements of 6.1% and 5.8%, respectively. This phenomenon can be attributed to the GLE module within SPM, which is specifically designed to capture local information. Consequently, SPM demonstrates a significant advantage in semantic segmentation tasks, where detailed local feature extraction is crucial.
Effectiveness on other Backbones
Visualization
Additionally, we employed the Effective Receptive Field (ERF) method (Luo et al. 2016) as a visualization tool to compare the changes in ERF before and after using SPM, as depicted in Fig. 5. It is readily observed that after integrating SPM, the size of the ERF not only increases significantly but also changes shape from a regular square to a radial decay pattern, which aligns more closely with biological vision. This pattern can be attributed to our carefully designed MSA and GLE modules, which together achieve an excellent balance between capturing long-range relationships and preserving local detail features. Consequently, this balance leads to improved performance in classification and downstream dense prediction tasks.
Ablation Study
The Effectiveness of GLE
To evaluate the effectiveness of the proposed GLE module, we conducted experiments by replacing GLE with two alternative methods, as shown in Tab. 5). Replacing GLE with a $2\times 2$ convolution or a $3\times 3$ average pooling layer results in a performance decrease of 2.3% and 2.4%, respectively.
Backbone Method #Params (M) Top-1 Acc. (%)
PVT-Tiny SPM (MSA + GLE) 14.0 79.5
GLE $\rightarrow$ $2\times 2$ Conv. 14.4 77.2 (-2.3)
GLE $\rightarrow$ $3\times 3$ AvgPool 12.8 77.1 (-2.4)
Table 5: Comparison between GLE and conventional pooling methods.
The Effectiveness of Guide Token
@@
We conducted comparative experiments to evaluate the effectiveness of the proposed $[Guide]$ token against two mainstream methods: the $[CLS]$ token (Dosovitskiy et al. 2020) and global average pooling (GAP) (Chu et al. 2021), as presented in Tab. 6. The results demonstrate that the $[Guide]$ token improves model performance by approximately 1.7% compared to these methods, without significantly increasing the number of parameters.
Backbone Method #Params (M) Top-1 Acc. (%)
PVT-Tiny GAP (Chu et al. 2021) 14.0 77.7 (-1.8)
Class token (Dosovitskiy et al. 2020) 14.0 77.9 (-1.6)
@@
Selection of GTG’s Kernel Size
To determine the optimal kernel size for GTG, we conducted a series of performance comparison experiments with different kernel sizes. Ultimately, we set the kernel size of GTG to 7, which achieved the best performance without significantly increasing the number of parameters. This indicates that information outside the local window positively influences the attention results within the window, highlighting the effectiveness of the $[Guide]$ token.
Backbone Kernel Size #Params (M) Top-1 Acc. (%)
PVT-Tiny $3\times 3$ 14.0 78.86
$5\times 5$ 14.0 79.43
$\bm{7\times 7}$ 14.0 79.47
$9\times 9$ 14.1 79.46
Table 7: Performance comparison of GTG with different kernel sizes.
Gradually Applying SPM
Different stages represent varying levels of semantic information. To validate the generalizability of the SPM framework across different levels of semantic information, we conducted experiments by gradually replacing the original patch merging with SPM (see Tab. 8). From the second and third rows, we observe that SPM enhances network performance by 1.3% with low-level features (stage 1) and by 1.6% with high-level features (stage 3), respectively. Furthermore, from the second, fourth, and fifth rows, it is evident that progressively replacing the original patch merging with SPM linearly improves performance by 1.3%, 3.0%, and 4.4%, respectively.
Backbone Stage1 Stage2 Stage3 #Params (M) Top-1 Acc. (%)
@@
Conclusion
In this work, we introduced the Stepwise Patch Merging framework, inspired by the brain’s ability to integrate global and local information for comprehensive visual understanding. The proposed SPM framework, comprising Multi-Scale Aggregation and Guided Local Enhancement modules, demonstrates significant improvements in various computer vision tasks, including classification, detection, and segmentation. Through extensive experiments on ImageNet-1K and COCO benchmarks, as well as on ADE20K for semantic segmentation, we showed that SPM consistently enhances the performance of different backbone models. The robustness of SPM to different input sizes and its effective generalization to other Transformer architectures further underscore its versatility and potential as a powerful enhancement technique. Future work will explore the application of SPM in more diverse domains and its integration with other state-of-the-art models to further elevate its impact on the field of computer vision.

### v2

@@ 4.2 Object Detection on COCO
Object detection and instance segmentation experiments were conducted on the challenging COCO benchmark Lin et al. (2014). All models were trained on the training set comprising 118k images and evaluated on the validation set with 5k images. We validated the effectiveness of different backbones using Mask R-CNN He et al. (2017). Before training, the weights pre-trained on ImageNet-1K were used to initialize the backbone, and the newly added layers were initialized using the Xavier initialization method Glorot and Bengio (2010). Our models were trained with a batch size of 16 on 8 NVIDIA A100 GPUs and optimized using the AdamW optimizer Loshchilov and Hutter (2017) with an initial learning rate of $1\times 10^{-4}$.
Result.
@@
As shown in Tab. 2, incorporating the SPM framework into the PVT resulted in significant improvements of 4.1%, 2.6%, and 1.3% in the Tiny, Small, and Medium models, respectively, for the object detection task. Models integrated with SPM show marked improvements in detecting medium-sized and large objects. This enhancement is attributed to the original patch merging’s relatively singular and small receptive fields, whereas the MSA module integrates features with diverse receptive fields, enabling the model to more accurately capture long-range dependencies. Moreover, there is a significant improvement in detecting small objects. Although larger models are typically better at modeling global relationships, the disruption of local information may hinder small objects from establishing complete semantic information, leading to missed detections. The GLE module addresses this by enhancing the perception of local discriminative information, resulting in consistent improvements in the detection performance of small objects across models of different sizes.
4.3 Semantic Segmentation on ADE20K
@@
ResNet18 He et al. (2016) 15.5 32.9
PVT-Tiny Wang et al. (2021) 17.0 35.7
PVT-Tiny (HAFA) Chen et al. (2023) 18.7 40.1
@@
ResNet50 He et al. (2016) 28.5 36.7
PVT-Small Wang et al. (2021) 28.2 39.8
PVT-Small (HAFA) Chen et al. (2023) 29.9 43.8
@@
ResNet101 He et al. (2016) 47.5 38.8
ResNeXt101-32$\times$4d Xie et al. (2017) 47.1 39.7
PVT-Medium Wang et al. (2021) 48.0 41.6
@@
Setting.
The ADE20K dataset Zhou et al. (2017) is a widely utilized benchmark for semantic segmentation, comprising 150 categories with 20,210 images for training, 2,000 images for validation, and 3,352 images for testing. All the methods compared were evaluated using the Semantic FPN framework Kirillov et al. (2019). The backbone network of our method was initialized with the pre-trained ImageNet-1k model, and the newly added layers were initialized using the Xavier initialization method. The initial learning rate was set to 0.0001, and the model was optimized using the AdamW optimizer. We trained our models for 40,000 iterations with a batch size of 16 on eight NVIDIA A100 GPUs. The learning rate followed a polynomial decay schedule with a power of 0.9. During training, images were randomly resized and cropped to $512\times 512$ pixels. For testing, images were rescaled to have a shorter side of 512 pixels.
Result.
As shown in Tab. 3, the integration of the SPM framework led to a significant enhancement in the semantic segmentation task. Specifically, the performance of the Tiny, Small, and Large models improved by 5.8%, 6.1%, and 3.7%, respectively. Interestingly, in both classification and detection tasks, the relative improvement brought by SPM compared to the base model gradually decreases as the model size increases. However, on the ADE20K dataset, the performance gain of PVT-Small exceeds that of PVT-Tiny, with improvements of 6.1% and 5.8%, respectively.
Figure 4: Visualization of the attention map includes the original images (the 1st column), the visualization of the attention map for each head after using SPM (the 2nd and 3rd columns), and without using SPM (the 4th and 5th columns). The red five-pointed star in the original images represents the position of the query.
4.4 Effectiveness on Other Backbones
4.5 Visualization
As shown in Fig. 5 (a), MSA enhances fine-grained features and moderate global context benefits detail generation, while excessive global aggregation leads to performance degradation (see Tab. 7).
Additionally, we employed the Effective Receptive Field (ERF) method Luo et al. (2016) as a visualization tool to compare the changes in ERF before and after using SPM, as depicted in Fig. 5. It is readily observed that after integrating SPM, the size of the ERF not only increases significantly but also changes shape from a regular square to a radial decay pattern, which aligns more closely with biological vision.
Figure 5: (a) Visualizations before/after MSA and GLE with increasing GTG kernel size (3, 5 and 7); (b) Visualization comparison of the Effective Receptive Field of PVT-Tiny before (the 1st row) and after applying our SPM (the 2nd row), using the outputs from the three stages separately. Each ERF image is generated by averaging over 5000 $256\times 256$ images from ImageNet-1K validation set.
@@ 4.5 Visualization
ICCV 2021 Swin-T Liu et al. (2021) 29.0 81.3
Swin-T (HAFA) Liu et al. (2021) 29.1 81.7
@@
CVPR 2023 NAT-Mini Hassani et al. (2023) 20.0 81.8
@@
5 Ablation Study
All experiments were conducted with the random seed fixed at 0, without tuning for desired outcomes.
5.1 The Effectiveness of MSA, GLE and GTG
As presented in Tab. 5, replacing GLE with a $3\times 3$ AvgPool results in a 2.4% performance decrease, but still outperforms the baseline by 2.0%, with nearly identical parameter count and computational complexity.
Although the DWConv in the GTG module introduces a small number of learnable parameters, it has been experimentally proven to be highly effective. For instance, replacing GTG with an AvgPool of the same kernel size results in a 1.5% performance drop.
Moreover, PVT-Tiny (SPM) achieves accuracy similar to PVT-Small while using fewer FLOPs and parameters, suggesting that SPM’s advantage comes from structural design rather than resource scaling.
Backbone Impl. of Patch Merging #P (M) #F (G) Top-1 Acc. (%)
PVT-Tiny SPM (MSA + GLE) 13.6 2.7 79.5
MSA $\rightarrow$ $3\times 3$ Conv. 15.1 3.3 78.8 (-0.7)
GLE $\rightarrow$ $2\times 2$ Conv. 14.0 2.2 77.2 (-2.3)
GLE $\rightarrow$ $3\times 3$ AvgPool 12.4 2.0 77.1 (-2.4)
GTG $\rightarrow$ $7\times 7$ AvgPool 13.6 2.7 78.0 (-1.5)
$2\times 2$ Conv. 12.8 1.9 75.1
PVT-Small $2\times 2$ Conv. 24.1 3.7 79.8
Table 5: Classification results on ImageNet-1K with different patch merging strategies. “#P” and “#F” are calculated by the thop package.
As shown in Fig. 6 (a), the additional 0.8 GFLOPs overhead mainly results from the linear projection in GLE, which fuses $[Guide]$ and regular tokens via self-attention. Initially, we experimented with replacing self-attention with dynamic weighting (SKNet Li et al. (2019)), which reduced FLOPs but caused a slight drop in performance. This highlights the trade-off between efficiency and accuracy, and supports our choice of maintaining attention in GLE for better feature fusion. We are also exploring more efficient guiding mechanisms for future work.
5.2 The Effectiveness of Guide Token
@@
We conducted comparative experiments to evaluate the effectiveness of the proposed $[Guide]$ token against two mainstream methods: the class token Dosovitskiy et al. (2020) and global average pooling (GAP) Chu et al. (2021), as presented in Tab. 6. The results demonstrate that the $[Guide]$ token improves model performance by approximately 1.7% compared to these methods, without significantly increasing the number of parameters.
Backbone Method #P (M) Top-1 Acc. (%)
PVT-Tiny GAP 14.0 77.7 (-1.8)
Class token 14.0 77.9 (-1.6)
@@
5.3 The Kernel Size of GTG
To determine the optimal kernel size for GTG, we conducted a series of performance comparison experiments with different kernel sizes, as shown in Tab. 7.
Backbone Kernel Size #P (M) Top-1 Acc. (%)
PVT-Tiny $1\times 1$ 13.6 78.17
$3\times 3$ 13.6 78.86
$5\times 5$ 13.6 79.43
$\bm{7\times 7}$ 13.6 79.47
$9\times 9$ 13.7 79.46
$31\times 31$ 14.5 79.34
Table 7: Performance comparison of GTG with different kernel sizes. “#P” is calculated by the thop package.
To assess optimization stability, we experimented with kernel sizes at extremes (1 and 31) and monitored training dynamics. As shown in Fig. 6 (b), training remains stable and converges faster than the baseline (PVT-Tiny), indicating that our $[Guide]$ token design does not introduce instability. Additionally, we had initially added residual connections to $[Guide]$ tokens, but found that removing them slightly improved performance, and the training remained stable throughout.
Figure 6: (a) FLOPs distribution across modules in SPM. (b) Stable convergence even with extreme kernel sizes (1 and 31).
5.4 Gradually Applying SPM
From Tab. 8, we conclude that 1) SPM is robust to varying input sizes containing different levels of semantic content, consistently improving performance (2nd and 3rd rows), and 2) repeated application of SPM leads to a healthy linear increase in performance (2nd, 4th and 5th rows).
Backbone Stage1 Stage2 Stage3 #P (M) Top-1 Acc. (%)
@@ 5.4 Gradually Applying SPM
6 Future Work
The current SPM involves many hyperparameters that need to be considered; future work will explore a more concise and computationally efficient stepwise structure to further enhance HVTs performance, and experiments demonstrate the effectiveness of the global-guide mechanism, which will be analyzed in greater depth.
7 Conclusion
In this work, we introduced the Stepwise Patch Merging, inspired by the brain’s ability to integrate global and local information for comprehensive visual understanding. The proposed SPM framework, comprising Multi-Scale Aggregation and Guided Local Enhancement modules, demonstrates significant improvements in various computer vision tasks, including classification, detection, and segmentation. Meanwhile, we showed that SPM consistently enhances the performance of different backbone models. The robustness of SPM to different input sizes and its effective generalization further underscore its versatility and potential as a powerful enhancement technique.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 62406325).